OpenStudy (anonymous):

what is the area of the parallelogram 40, 68 and 150 degrees

OpenStudy (calculusxy):

The area of a parallelogram is given with the formulae of A=b*h.

OpenStudy (anonymous):

I Know but I don't know how to resolve the problem

OpenStudy (anonymous):

Use the formula , Area of parallelogram = ab sin C...that is two adjacent sides multiplied by the sine of the included angle. So, Area = 40(68) sin (150) = 40(68)(1/2) 1,360

OpenStudy (anonymous):

Honestly, I cannot understand what you are asking me. Perhaps rephrase what you are asking.

OpenStudy (anonymous):

How do you get 1,360

OpenStudy (anonymous):

40 x 68 x sin 150 degrees

OpenStudy (anonymous):

put it on your calculator

OpenStudy (anonymous):

got it?

OpenStudy (anonymous):

thank you so much! now I'm ready for the SPAT

OpenStudy (anonymous):

Good Luck.

OpenStudy (anonymous):

You mean PSAT

OpenStudy (anonymous):

Thanks